Can someone show me how to solve this problem?Question 8 On July 31, 2017, Blue Company engaged Minsk Tooling Company to construct a special purpose piece of factory machi
Answer #1
Interest revenue $2550
Weighted average accumulated expenditure $48600
Avoidable interest $5832
Interest capitalized $5832

Explanation:

Interest revenue = $102000 × 10% × 3/12 = $2550

Date expenditure (a) capitalization period (b) weight (c = b/12) weighted expenditure (a×c)
July 31 $194400 3 0.25 $48600
Nov. 01 $102000 0 0 0
$296400 $48600

Avoidable interest = $48600 × 12% = $5832

Interest capitalized = $5832

.

.

PLEASE DO GIVE A THUMBS UP